# 3D Gaussian Parametric Head Model

### v1

@@ 3.2 Model Representation
$\{S,Q_{can},A\}=\{\boldsymbol{S_{0}},\boldsymbol{Q_{0}},\boldsymbol{A_{0}}\}+\boldsymbol{f_{att}}(H,\boldsymbol{z^{exp}}).$ (4)
Following this, we utilize the estimated head pose parameters $R$ and $T$, obtained during data preprocessing, to transform the canonical space variables $X_{can}$ and $Q_{can}$ into the world space: $X=R\cdot X_{can}+T,\ Q=R\cdot Q_{can}$. For model rendering, we leverage differentiable rendering [21] and neural rendering techniques to generate images. The generated 3D Gaussian parameters, which include $\{X,C,S,Q,A\}$, are conditioned by the identity code $\boldsymbol{z^{id}}$ and expression code $\boldsymbol{z^{exp}}$. Finally, we input this feature map into the same upsampling network $\boldsymbol{\Psi}(\cdot)$ of the guiding geometry model to generate a 512-resolution RGB image.
In the 3D Gaussian Parametric Head Model, we leverage the previously trained guiding geometry model to initialize our variables and networks, rather than initiating them randomly and training from scratch. Specifically, we initialize the Gaussian positions $\boldsymbol{X_{0}}$ using the vertex positions of the mean mesh $V_{0}$. Meanwhile, we generate the per-vertex feature $\Gamma$ from $\boldsymbol{f_{mean}}(x)$ at the beginning and bind it to the points as an optimizable variable $\boldsymbol{\Gamma_{0}}$ as described above. Additionally, all identity codes $\boldsymbol{z^{id}}$, expression codes $\boldsymbol{z^{exp}}$, and the networks $\{\boldsymbol{f_{inj}}(\cdot),\boldsymbol{f_{id}}(\cdot),\boldsymbol{f_{exp}}(\cdot),\boldsymbol{f_{col}}(\cdot),\boldsymbol{\Psi}(\cdot)\}$ are directly inherited from the guiding geometry model. Note that, the attribute MLP $\boldsymbol{f_{att}}(\cdot)$ is a newly introduced network, hence it is initialized randomly. Finally, the overall mean values of the Gaussian attributes $\{\boldsymbol{S_{0}},\boldsymbol{Q_{0}},\boldsymbol{A_{0}}\}$ are initialized following the original 3D Gaussian Splatting [21].
@@ 3.3 Loss Functions
To ensure the accurate convergence of the model, we employ various loss functions as constraints, including the basic photometric loss and silhouette loss, to enforce consistency with ground truth of both the rendered high-resolution images $I_{hr}$ and the rendered masks $M$:
$\mathcal{L}_{hr}=||I_{hr}-I_{gt}||_{1},\ \mathcal{L}_{sil}=IOU(M,M_{gt}),$ (5)
with $I_{gt}$ representing the ground truth RGB images, $M_{gt}$ representing the ground truth masks. We further encourage the first three channels of the low-resolution feature map $I_{lr}$ to closely match the ground-truth RGB image $I_{gt}$ by introducing an $L_{1}$ loss:
$\mathcal{L}_{lr}=||I_{lr}-I_{gt}||_{1}.$ (6)
@@
$\mathcal{L}_{lmk}=||\boldsymbol{P}-\boldsymbol{P}_{gt}||_{2},$ (7)
with $\boldsymbol{P}_{gt}$ denoting the ground truth 3D landmarks, which are estimated by fitting a BFM model to the training data during preprocessing.
Moreover, to guarantee the decoupling of identity and expression deformations learned by the model and minimize redundancy, we introduce the following regularization loss function that aims to minimize the magnitude of both deformations:
$\mathcal{L}_{reg}=||\delta V_{id}||_{2}+||\delta V_{exp}||_{2}.$ (8)
@@
$\begin{split}\mathcal{L}=\mathcal{L}_{hr}+\lambda_{sil}\mathcal{L}_{sil}+\lambda_{lr}\mathcal{L}_{lr}+\lambda_{lmk}\mathcal{L}_{lmk}+\lambda_{reg}\mathcal{L}_{reg}+\lambda_{lap}\mathcal{L}_{lap}\end{split}$ (9)
@@
with all the $\lambda$ denoting the weights of each term. In practice, we set $\lambda_{sil}=0.1$, $\lambda_{lr}=0.1$, $\lambda_{lmk}=0.1$, $\lambda_{reg}=0.001$ and $\lambda_{lap}=100$. During training, we jointly optimize the bolded variables above: $\boldsymbol{z^{id}}$, $\boldsymbol{z^{exp}}$, $\boldsymbol{f_{inj}}(\cdot)$, $\boldsymbol{f_{mean}}(\cdot)$, $\boldsymbol{f_{id}}(\cdot)$, $\boldsymbol{f_{exp}}(\cdot)$, $\boldsymbol{f_{col}}(\cdot)$, $\boldsymbol{\Psi}(\cdot)$, $\boldsymbol{P_{0}}$}. Notably, the defined canonical 3D landmarks $\boldsymbol{P_{0}}$ are initialized by computing the average of the estimated 3D landmarks from the training dataset.
@@
During the training stage of the 3D Gaussian Parametric Head Model, we also calculate the perceptual loss [60] to encourage the model to learn more high-frequency details $\mathcal{L}_{vgg}=VGG(I_{hr},I_{gt})$. Similar to training the guiding geometry model, we enforce the first three channels of the feature map to be RGB channels as Eqn. 6, introduce landmarks guidance terms as Eqn. 7 and the regular term for the displacement of points as Eqn. 8. Consequently, the overall loss function can be formulated as:
$\begin{split}\mathcal{L}=\mathcal{L}_{hr}+\lambda_{vgg}\mathcal{L}_{vgg}+\lambda_{lr}\mathcal{L}_{lr}+\lambda_{lmk}\mathcal{L}_{lmk}+\lambda_{reg}\mathcal{L}_{reg}\end{split}$ (10)
@@
with the weights $\lambda_{vgg}=0.1$, $\lambda_{lr}=0.1$, $\lambda_{lmk}=0.1$ and $\lambda_{reg}=0.001$. In this training stage, we also jointly optimize all the bolded variables and networks mentioned above, including the overall mean positions and attributes of the Gaussians and the 3D landmarks: $\boldsymbol{z^{id}}$, $\boldsymbol{z^{exp}}$, $\boldsymbol{f_{inj}}(\cdot)$, $\boldsymbol{f_{id}}(\cdot)$, $\boldsymbol{f_{exp}}(\cdot)$, $\boldsymbol{f_{col}}(\cdot)$, $\boldsymbol{f_{att}}(\cdot)$, $\boldsymbol{\Psi}(\cdot)$, $\boldsymbol{X_{0}}$, $\boldsymbol{\Gamma_{0}}$, $\boldsymbol{S_{0}}$, $\boldsymbol{Q_{0}}$, $\boldsymbol{A_{0}}$, $\boldsymbol{P_{0}}$}.
3.4 Inference Details
Image-based Fitting. When a single RGB portrait image is input, we first align the image according to the processing rules of the training set. Subsequently, we employ gradient descent to fit the image rendered by the 3D Gaussian Parametric Head Model to this input image using the photometric loss $\mathcal{L}_{lr}$ and $\mathcal{L}_{hr}$ defined in Eqn. 10. This process helps regress the identity code $\boldsymbol{z^{id}}$ and expression code $\boldsymbol{z^{exp}}$. We just optimize for 200 iterations with learning rate $1\times 10^{-3}$ for both latent codes. Following this, we fix the latent codes $\boldsymbol{z^{id}}$ and $\boldsymbol{z^{exp}}$, such that the variables $H$, $X_{can}$ are also fixed. We further optimize the color MLP $\boldsymbol{f_{col}}(\cdot)$ and the canonical positions $X_{can}$ which represent the geometry of the current specific subject, using the same loss function. In this step, we only optimize for 100 iterations with learning rate $1\times 10^{-4}$ for both $\boldsymbol{f_{col}}(\cdot)$ and $X_{can}$. This optimization process aims to add some details that cannot be recovered by the trained model itself, ultimately resulting in the reconstructed head model. The entire process has a total of 300 iterations and takes only 30 seconds.
Expression Editing. Given a source portrait image providing the subject whose expression is to be edited and a target portrait image providing the target expression. We first obtain the head model of the source subject through optimization as the above-mentioned Image-based Fitting strategy. Then for the target portrait image, we also obtain the head model and corresponding expression code in the same way. Finally, we input the target expression code to the head model of the source subject, so that the expression of the source subject can be edited to the target one.

### v2

@@ 3.2 Model Representation
$\{S,Q_{can},A\}=\{\boldsymbol{S_{0}},\boldsymbol{Q_{0}},\boldsymbol{A_{0}}\}+\boldsymbol{f_{att}}(H).$ (11)
Also at this stage, we do not consider expression-related Gaussian attributes changes as color changes mentioned above.
Following this, we utilize the estimated head pose parameters $R$ and $T$, obtained during data preprocessing, to transform the canonical space variables $X_{can}$ and $Q_{can}$ into the world space: $X=R\cdot X_{can}+T,\ Q=R\cdot Q_{can}$. For model rendering, we leverage differentiable rendering [22] and neural rendering techniques to generate images. The generated 3D Gaussian parameters, which include $\{X,C,S,Q,A\}$, are conditioned by the identity code $\boldsymbol{z^{id}}$, the face image $I_{face}$ and the non-face image $I_{nonface}$. Finally, we input this feature map into the same refine network $\boldsymbol{\Psi}(\cdot)$ of the guiding geometry model to generate a 512-resolution RGB image.
In the 3D Gaussian Parametric Head Model, we leverage the previously trained guiding geometry model to initialize our variables and networks, rather than initiating them randomly and training from scratch. Specifically, we initialize the Gaussian positions $\boldsymbol{X_{0}}$ using the vertex positions of the mean mesh $V_{0}$. Meanwhile, we generate the per-vertex feature $\Gamma$ from $\boldsymbol{f_{mean}}(x)$ at the beginning and bind it to the points as an optimizable variable $\boldsymbol{\Gamma_{0}}$ as described above. Additionally, all identity codes $\boldsymbol{z^{id}}$, 3D landmarks $\boldsymbol{P}_{0}$ and the networks $\boldsymbol{E}_{exp}(\cdot)$, $\boldsymbol{E}_{nf}(\cdot)$, $\{\boldsymbol{f_{inj}}(\cdot),\boldsymbol{f_{id}}(\cdot),\boldsymbol{f_{exp}}(\cdot),\boldsymbol{f_{nf}}(\cdot),\boldsymbol{f_{col}}(\cdot),\boldsymbol{\Psi}(\cdot)\}$ are directly inherited from the guiding geometry model. Note that, the attribute MLP $\boldsymbol{f_{att}}(\cdot)$ is a newly introduced network, hence it is initialized randomly. Finally, the overall mean values of the Gaussian attributes $\{\boldsymbol{S_{0}},\boldsymbol{Q_{0}},\boldsymbol{A_{0}}\}$ are initialized following the original 3D Gaussian Spatting [22].
@@ 3.3 Loss Functions
To ensure the accurate convergence of the model, we employ various loss functions as constraints, including the basic photometric loss and silhouette loss, to enforce consistency with ground truth of both the rendered fine images $I_{fine}$ and the rendered masks $M$:
$\mathcal{L}_{fine}=||I_{fine}-I_{gt}||_{1},$ (12)
$\mathcal{L}_{sil}=IOU(M,M_{gt}),$ (13)
with $I_{gt}$ representing the ground truth RGB images, $M_{gt}$ representing the ground truth masks. We further encourage the first three channels of the feature map $I_{coarse}$ to closely match the ground-truth RGB image $I_{gt}$ by introducing an $L_{1}$ loss:
$\mathcal{L}_{coarse}=||I_{coarse}-I_{gt}||_{1}.$ (14)
@@
$\mathcal{L}_{lmk}=||\boldsymbol{P}-P_{gt}||_{2},$ (15)
with $P_{gt}$ denoting the ground truth 3D landmarks, which are estimated by fitting a BFM model to the training data during preprocessing.
Moreover, to guarantee the decoupling of identity deformations and motion deformations learned by the model and minimize redundancy, we introduce the following regularization loss function that aims to minimize the magnitude of motion deformations:
$\mathcal{L}_{reg}=||\delta V_{exp}||_{2}+||\delta V_{nf}||_{2}.$ (16)
@@
$\mathcal{L}=$ $\mathcal{L}_{fine}+\lambda_{sil}\mathcal{L}_{sil}+\lambda_{coarse}\mathcal{L}_{coarse}+$ (17)
$\lambda_{lmk}\mathcal{L}_{lmk}+\lambda_{reg}\mathcal{L}_{reg}+\lambda_{lap}\mathcal{L}_{lap}$ (18)
@@
with all the $\lambda$ denoting the weights of each term. In practice, we set $\lambda_{sil}=0.1$, $\lambda_{coarse}=0.1$, $\lambda_{lmk}=0.1$, $\lambda_{reg}=0.001$ and $\lambda_{lap}=100$. During training, we jointly optimize the bolded variables above: $\boldsymbol{z^{id}}$, $\boldsymbol{E}_{exp}(\cdot)$, $\boldsymbol{E}_{nf}(\cdot)$, $\boldsymbol{f_{inj}}(\cdot)$, $\boldsymbol{f_{mean}}(\cdot)$, $\boldsymbol{f_{id}}(\cdot)$, $\boldsymbol{f_{exp}}(\cdot)$, $\boldsymbol{f_{nf}}(\cdot)$, $\boldsymbol{f_{col}}(\cdot)$, $\boldsymbol{\Psi}(\cdot)$, $\boldsymbol{P_{0}}$}. Notably, the defined canonical 3D landmarks $\boldsymbol{P_{0}}$ are initialized by computing the average of the estimated 3D landmarks from the training dataset.
@@
During the training stage of the 3D Gaussian Parametric Head Model, we also calculate the perceptual loss [78] to encourage the model to learn more high-frequency details $\mathcal{L}_{vgg}=VGG(I_{fine},I_{gt})$. Similar to training the guiding geometry model, we enforce the first three channels of the feature map to be RGB channels as Eqn. 14, introduce landmarks guidance terms as Eqn. 15 and the regular term for the displacement of points as Eqn. 16. Consequently, the overall loss function can be formulated as:
$\mathcal{L}=$ $\mathcal{L}_{fine}+\lambda_{vgg}\mathcal{L}_{vgg}+\lambda_{coarse}\mathcal{L}_{coarse}+$ (19)
$\lambda_{lmk}\mathcal{L}_{lmk}+\lambda_{reg}\mathcal{L}_{reg}$ (20)
@@
with the weights $\lambda_{vgg}=0.1$, $\lambda_{coarse}=0.1$, $\lambda_{lmk}=0.1$ and $\lambda_{reg}=0.001$. In this training stage, we also jointly optimize all the bolded variables and networks mentioned above, including the overall mean positions and attributes of the Gaussians and the 3D landmarks: $\boldsymbol{z^{id}}$, $\boldsymbol{E}_{exp}(\cdot)$, $\boldsymbol{E}_{nf}(\cdot)$, $\boldsymbol{f_{inj}}(\cdot)$, $\boldsymbol{f_{id}}(\cdot)$, $\boldsymbol{f_{exp}}(\cdot)$, $\boldsymbol{f_{nf}}(\cdot)$, $\boldsymbol{f_{col}}(\cdot)$, $\boldsymbol{f_{att}}(\cdot)$, $\boldsymbol{\Psi}(\cdot)$, $\boldsymbol{X_{0}}$, $\boldsymbol{\Gamma_{0}}$, $\boldsymbol{S_{0}}$, $\boldsymbol{Q_{0}}$, $\boldsymbol{A_{0}}$, $\boldsymbol{P_{0}}$}.
3.4 Training Details
Before training starts, we first initialize the identity codes following NPHM [17]. For each different identity, we set a different identity code with a dimension of 512. In addition, we enforce a constraint to ensure that the norm of these codes remains below 1.
Besides, we experimentally found that although it is not necessary, jointly optimizing the head pose during the training process can eliminate some minor errors generated during the BFM calibration process, promoting the consistency of the model as the codes change. Consequently, we optimize all the head poses with a very small learning rate $1\times 10^{-5}$ throughout the entire training stage of the model.
For the face image in the training process, we randomly select the image of another view of the current frame, or the synthesized image by LivePortrait [30] with the same expression as the input.
Next, we give a brief description of the network structure. The each of the 2 encoders $\boldsymbol{E}_{exp}(\cdot)$, $\boldsymbol{E}_{nf}(\cdot)$ consists of 4 convolution blocks plus an MLP. Each block contains a convolutional layer with stride 1 and a convolutional layer with stride 2. For the refine network $\boldsymbol{\Psi}(\cdot)$, we adopt a very simple U-net [79] structure, with 2 convolutional layer for downsampling and 2 convolutional layers for upsampling. For the MLPs $\boldsymbol{f_{mean}}(\cdot)$, $\boldsymbol{f_{id}}(\cdot)$, $\boldsymbol{f_{exp}}(\cdot)$, $\boldsymbol{f_{nf}}(\cdot)$, $\boldsymbol{f_{col}}(\cdot)$, $\boldsymbol{f_{att}}(\cdot)$, we set the width to 512 with 4 hidden layers. And for the injection MLP $\boldsymbol{f_{inj}}(\cdot)$, we set the width to 512 with 8 hidden layers. The mesh and the Gaussians are both rendered as 512-resolution feature maps and transferred into 512-resolution RGB images through the refine network.
During training the guiding geometry model, we use 256-resolution tetrahedral grid for extracting the mesh via DMTet. For the optimization, we use an Adam [80] optimizer, and set the learning rate to $1\times 10^{-4}$ for the identity codes $\boldsymbol{z^{id}}$, $1\times 10^{-4}$ for all the networks $\boldsymbol{E}_{exp}(\cdot)$, $\boldsymbol{E}_{nf}(\cdot)$, $\boldsymbol{f_{inj}}(\cdot)$, $\boldsymbol{f_{mean}}(\cdot)$, $\boldsymbol{f_{id}}(\cdot)$, $\boldsymbol{f_{exp}}(\cdot)$, $\boldsymbol{f_{nf}}(\cdot)$, $\boldsymbol{f_{col}}(\cdot)$, $\boldsymbol{\Psi}(\cdot)$, and $1\times 10^{-4}$ for the 3D landmarks $\boldsymbol{P_{0}}$. We use a batch size of 8, with each batch containing 4 images of a specific expression from a given identity. Training the guiding geometry model requires 8 RTX4090 graphics cards and approximately 1 day.
While training the Gaussian model, we also use an Adam optimizer and set the learning rate: $1\times 10^{-5}$ for the identity codes $\boldsymbol{z^{id}}$, $1\times 10^{-4}$ for all the networks $\boldsymbol{E}_{exp}(\cdot)$, $\boldsymbol{E}_{nf}(\cdot)$, $\boldsymbol{f_{id}}(\cdot)$, $\boldsymbol{f_{exp}}(\cdot)$, $\boldsymbol{f_{nf}}(\cdot)$, $\boldsymbol{f_{col}}(\cdot)$, $\boldsymbol{f_{att}}(\cdot)$, $\boldsymbol{\Psi}(\cdot)$ and $1\times 10^{-4}$ for the 3D landmarks $\boldsymbol{P_{0}}$. For the mean Gaussian attributes, we set the learning rates as: $1\times 10^{-5}$ for the positions $\boldsymbol{X_{0}}$, $1\times 10^{-5}$ for the per-vertex feature $\boldsymbol{\Gamma_{0}}$, $3\times 10^{-5}$ for the scale $\boldsymbol{S_{0}}$, $1\times 10^{-5}$ for the rotation $\boldsymbol{Q_{0}}$ and $1\times 10^{-4}$ for the opacity $\boldsymbol{A_{0}}$. We use a batch size of 8, with each batch containing a single image. Following the training of the guiding geometry model, we transfer its parameters to the Gaussian model and continue training on 8 RTX4090 graphics cards for 7 days until convergence is achieved.
Figure 4: The pipeline of head avatar reconstruction from monocular videos. First, we optimize the identity code $\boldsymbol{z}_{id}$ to coarsely fit the GPHM model to the input video. Then we directly finetune the 3D Gaussian attributes and the motion-related networks in the GPHM for a fine-grained head avatar. The flame chart in the figure marks the parameters that need to be optimized.
3.5 Head Avatar from a Monocular Video
Once we have trained the Gaussian Parametric Head Model, one of the main applications is fast reconstruction for 3D head avatars from monocular videos, or even from few-shot or single-image inputs. Previous methods [3, 4, 5, 28, 29] trained their model from scratch and used the 3DMM expression coefficients as the driving signal. Therefore, when training the model, sufficient data is required to ensure generalization and reconstruction quality, and the subsequent driving process requires tracking for the 3DMM expression of the source actor. In contrast, we utilize the well-learned generalizable prior model for both appearance and expression to model head avatars. Specifically, we first fit the GPHM to the input video to obtain a rough 3D Gaussian model in a few iterations. Further, we just slightly finetune the 3D Gaussian parameters and the refine network to reconstruct a fine-grained Gaussian model. Our pipeline achieves faster reconstruction and better rendering quality while requiring less training data. In addition, as our framework leverages a pretrained expression encoder for end-to-end expression control, the head avatar can be animated directly by a face image or video, without tracking for 3DMM expression. In the experiment, we also verified that this expression control strategy shows better generalization ability.
GPHM Fitting At this stage, we first optimize the identity code to obtain a coarsely fitting GPHM model. Given a $N$ frame monocular frontal portrait video, a set of few-shot images, or even a single image input $\{{I_{n}}\},n\in\{1,...,N\}$, we first remove the background of the images, then detect 2D landmarks for extracting face images $\{{I_{face}^{n}}\},n\in\{1,...,N\}$, non-face images $\{{I_{nonface}^{n}}\},n\in\{1,...,N\}$ and estimating the head pose $\{{R_{n},T_{n}}\},n\in\{1,...,N\}$ in the same way as the training data preprocessing described in Section 3.1. Then, we randomly initialize a global identity code $\boldsymbol{z^{id}}$ and fit our GPHM to the input video by optimizing the identity code. Specifically, in each iteration, we sample one frame $n$ and input the identity code $\boldsymbol{z^{id}}$, the face image ${I_{face}^{n}}$ and the non-face image ${I_{nonface}^{n}}$ to the GPHM to generate 3D Gaussians $\{X_{n},C_{n},S_{n},Q_{n},A_{n}\}$ as described above. We then render the feature map $I_{F}$ through rasterization with the first three channels as the coarse RGB image $I_{coarse}$. Finally the feature map $I_{F}$ is transferred to fine image $I_{fine}$ through the refine network $\boldsymbol{\Psi}$. For the loss function, we use only photometric loss $\mathcal{L}_{coarse}$ and $\mathcal{L}_{fine}$ defined in Eqn. 19. We totally optimize the identity code $\boldsymbol{z^{id}}$ for 100 iterations with learning rate $1\times 10^{-3}$.
Finetune 3D Gaussians In this stage, we further optimize the Gaussian attributes to reconstruct a high-fidelity identity-specific head avatar. First, we calculate the Gaussians positions $\boldsymbol{X_{id}}=\boldsymbol{X_{0}}+\delta X_{id}$, the color and other Gaussian attributes $\boldsymbol{C_{id}}$, $\boldsymbol{S_{id}}$, $\boldsymbol{Q_{id}}$, $\boldsymbol{A_{id}}$ as Eqn. 10 and Eqn. 11 using the identity code obtained in the GPHM fitting stage and set them as optimizable variables. As the computations are completed, we no longer need the networks $\boldsymbol{f_{inj}}(\cdot)$, $\boldsymbol{f_{id}}(\cdot)$, $\boldsymbol{f_{col}}(\cdot)$, $\boldsymbol{f_{att}}(\cdot)$ and the identity code. So they are subsequently discarded for saving computational overhead. Then, we consider modeling some identity-specific dynamic details deriving from the color and Gaussian attributes varying with the expression. Referring to the approach in Gaussian Head Avatar [54], we introduce an additional tiny MLP $\boldsymbol{f_{dyn}}(\cdot)$, which takes the facial expression code $z^{exp}$ and the per-Gaussian feature $\gamma$ as input and predicts the offset of the variables $\{\delta C,\delta S,\delta Q,\delta A\}$. Finally, the optimization process is the same as the GPHM Fitting stage. In each iteration, one frame $n$ is sampled, the face image ${I_{face}^{n}}$ and the non-face image ${I_{nonface}^{n}}$ are input to the model and render the feature map $I_{F}$ which is transferred to the fine image $I_{fine}$ later. Finally, we construct the same loss function as Eqn. 19. Typically, during finetuning, we optimize the identity-specific Gaussians $\boldsymbol{X_{id}}$, $\boldsymbol{C_{id}}$, $\boldsymbol{S_{id}}$, $\boldsymbol{Q_{id}}$, $\boldsymbol{A_{id}}$ with learning rate $1\times 10^{-3}$ and the networks $\boldsymbol{f_{exp}}(\cdot)$, $\boldsymbol{\Psi}(\cdot)$, $\boldsymbol{f_{dyn}}(\cdot)$ with learning rate $3\times 10^{-4}$ for 2000 iterations. But in the case of few-shot input, the optimization of the facial expression MLP $\boldsymbol{f_{exp}}(\cdot)$ and refine network $\boldsymbol{\Psi}(\cdot)$ is optionally turned off to prevent overfitting and reduce the number of iterations appropriately according to the number of input images.
Reenactment Once the identity-specific head avatar is reconstructed, we can use another portrait video for reenactment. Given a frame of the driving video, we first estimate the head pose and extract face image and non-face image as training data preprocessing 3.1. Then, we feed the face image to the facial expression encoder to generate the facial expression code for expression controlling. Similarly, we also feed the non-face image to the non-face motion encoder to generate non-face motion code to control the neck and shoulder motion when turning the head. Note, optionally the control of the neck and below could be ignored by fixing the non-face motion code as a sample in the training set. Finally, we input two motion codes to the finetuned model to generate 3D Gaussians, which is rendered into images.